# Hybrid Wi-Fi/PDR Indoor Localization with Fingerprint Matching

### v1

@@ 1 Introduction
With the advancement of wireless communication technology and the popularity of smartphones, location-based service (LBS) has attracted much attention [16]. LBS refers to software services that use geographic data and information to provide services or information to users, typically used in navigation software, social networking services, location-based advertising, tracking systems, mobile commerce, and personalized weather services [36]. Its implementation is based on determining the user’s location information, namely positioning technology. Traditional positioning technologies use satellites for positioning, such as the Global Positioning System (GPS) and China’s Beidou Satellite Navigation System [28]. Satellite-based positioning technology is mature, accurate, stable, and has a good positioning effect in the outdoor environment. However, because of the low power of the satellite signal, it is difficult to penetrate the walls of the building. Coupled with its own scattering, reflection, multipath effect, etc., the signal has been severely attenuated in the indoor environment [38]. Due to the influence of the complex indoor environment, the positioning effect has been far from meeting the needs of people. In addition, indoor positioning requires higher positioning accuracy, so the development of indoor positioning systems has become a research hotspot in the Internet of Things.
@@ 2 Related Work
The main indoor positioning techniques can be divided into three categories: computer vision-based, sensor-based, and communication-based positioning [18]. Computer vision-based systems employ omnidirectional cameras, 3D cameras, or built-in smartphone cameras to extract information about the indoor environment. They then use various image processing algorithms for feature extraction and matching [26]. Although this method boasts high accuracy, it also has significant financial implications and is susceptible to external factors that can disrupt signals. Sensor-based methods principally refer to Pedestrian Dead Reckoning (PDR), which uses data from accelerometers, gyroscopes, magnetometers, etc., and can estimate the user’s position based on past positions [21]. However, this method is frequently integrated with other positioning techniques due to the accumulation of drift-related errors over time and the requirement for initial calibration [24].
@@ 3 Methodology
Location fingerprint technology is an implementation scheme of indoor positioning technology based on Wi-Fi [51]. Despite the advantages of Wi-Fi, including its extensive deployment, ease of promotion and low cost, its use as a communication signal is generally limited to a single antenna with a small bandwidth, which hinders its compatibility with traditional positioning and ranging methods. The advent of location fingerprint technology has effectively addressed this issue [8]. Location fingerprinting is defined as the process of associating a location with a unique identifier, known as a ”fingerprint”. The location is defined as the coordinate of the positioning point, and the fingerprint is defined as the feature information of the positioning point. As demonstrated in Figure 1, the process of location fingerprint localization is comprised of two distinct stages: offline training and online localization [50]. In the offline phase, location fingerprints are collected and a wireless map is established (see Section 3.1). In the online phase, fingerprint samples are best matched with fingerprints in the wireless map to obtain location coordinates (see Section 3.2). Furthermore, the system integrates a PDR algorithm to achieve the trajectory tracking function (see Section 3.3), based on the initial position obtained through the positioning function. Finally, Section 3.4 provides a comprehensive overview of the system architecture.
@@ 3.1 Location Fingerprint Database Construction
In principle, any information feature that can identify a location can be used to create a location fingerprint [30]. Such features include the multipath structure of the Wi-Fi signal at a given location, the monitoring of an access station or a public mobile communication base station at a location, the RSS from the base station that can be monitored at a location, the round-trip time of the signals during communication at a location, and the delay [15], etc. The location fingerprint employed in this system is the RSS (Radio Signal Strength) signal, which is also one of the most commonly used signal features. RSS is a special term for RF (Radio Frequency) signals, which refers to the power output of a transmitter received by a reference antenna at a certain distance from the transmitting antenna [35]. The strength of the received signal is used to determine the distance between the point of the signal and the point of reception. The calculation formula for RSS is shown in Equation 1 [40]:
@@ 3.4 System Architecture
The system utilises a MySQL database for the purpose of data storage. The data to be stored comprises the location coordinates (X and Y) and the signal strength of the Wi-Fi deployment point [39]. The data types are of the float variety, and none of them permit null values. The design is illustrated in Table 1.
@@ 5 Conclusion
This paper preliminarily explores the classification of indoor position technology and the implementation scheme based on location fingerprint technology. Common indoor localization techniques include computer vision-based localization, sensor-based localization, and localization based on different communication technologies such as RFID, Wi-Fi, Bluetooth, visible light communication, and ultra-wideband. Location fingerprinting has emerged as a prominent technique in this field, using widely deployed Wi-Fi networks. This study proposes a pragmatic hybrid indoor positioning system that integrates Wi-Fi fingerprinting and pedestrian dead reckoning, providing a cost-effective solution that optimizes accuracy, real-time performance, and deployment complexity. By using existing Wi-Fi infrastructure and smartphone sensors, the system can perform data acquisition, localization, and trajectory tracking functions without additional hardware. The proposed dynamic weighted WKNN algorithm enhances traditional fingerprint matching by incorporating distance-based weighting, while sensor fusion with PDR effectively achieves real-time trajectory tracking. The experimental results obtained from field tests have confirmed the feasibility of the system, with stable localization accuracy and low latency.

### v2

@@ 1 Introduction
With the advancement of wireless communication technology and the popularity of smartphones, location-based service (LBS) has attracted much attention [26, 68, 69]. LBS refers to software services that use geographic data and information to provide services or information to users, typically used in navigation software, social networking services, location-based advertising, tracking systems, mobile commerce, and personalized weather services [50, 1]. Its implementation is based on determining the user’s location information, namely positioning technology. Traditional positioning technologies use satellites for positioning, such as the Global Positioning System (GPS) and China’s Beidou Satellite Navigation System [41, 2]. Satellite-based positioning technology is mature, accurate, stable, and has a good positioning effect in the outdoor environment. However, because of the low power of the satellite signal, it is difficult to penetrate the walls of the building. Coupled with its own scattering, reflection, multipath effect, etc., the signal has been severely attenuated in the indoor environment [52]. Due to the influence of the complex indoor environment, the positioning effect has been far from meeting the needs of people. In addition, indoor positioning requires higher positioning accuracy, so the development of indoor positioning systems has become a research hotspot in the Internet of Things.
@@ 2 Related Work
The main indoor positioning techniques can be divided into three categories: computer vision-based, sensor-based, and communication-based positioning [29, 3]. Computer vision-based systems employ omnidirectional cameras, 3D cameras, or built-in smartphone cameras to extract information about the indoor environment. They then use various image processing algorithms for feature extraction and matching [40, 9]. Although this method boasts high accuracy, it also has significant financial implications and is susceptible to external factors that can disrupt signals. Sensor-based methods principally refer to Pedestrian Dead Reckoning (PDR), which uses data from accelerometers, gyroscopes, magnetometers, etc., and can estimate the user’s position based on past positions [33, 10]. However, this method is frequently integrated with other positioning techniques due to the accumulation of drift-related errors over time and the requirement for initial calibration [37, 62].
@@ 3 Methodology
Location fingerprint technology is an implementation scheme of indoor positioning technology based on Wi-Fi [71, 57]. Despite the advantages of Wi-Fi, including its extensive deployment, ease of promotion and low cost, its use as a communication signal is generally limited to a single antenna with a small bandwidth, which hinders its compatibility with traditional positioning and ranging methods. The advent of location fingerprint technology has effectively addressed this issue [13, 17]. Location fingerprinting is defined as the process of associating a location with a unique identifier, known as a ”fingerprint”. The location is defined as the coordinate of the positioning point, and the fingerprint is defined as the feature information of the positioning point. As demonstrated in Figure 1, the process of location fingerprint localization is comprised of two distinct stages: offline training and online localization [70, 18]. In the offline phase, location fingerprints are collected and a wireless map is established (see Section 3.1). In the online phase, fingerprint samples are best matched with fingerprints in the wireless map to obtain location coordinates (see Section 3.2). Furthermore, the system integrates a PDR algorithm to achieve the trajectory tracking function (see Section 3.3), based on the initial position obtained through the positioning function. Finally, Section 3.4 provides a comprehensive overview of the system architecture.
@@ 3.1 Location Fingerprint Database Construction
In principle, any information feature that can identify a location can be used to create a location fingerprint [43, 19]. Such features include the multipath structure of the Wi-Fi signal at a given location, the monitoring of an access station or a public mobile communication base station at a location, the RSS from the base station that can be monitored at a location, the round-trip time of the signals during communication at a location, and the delay [25, 20], etc. The location fingerprint employed in this system is the RSS (Radio Signal Strength) signal, which is also one of the most commonly used signal features. RSS is a special term for RF (Radio Frequency) signals, which refers to the power output of a transmitter received by a reference antenna at a certain distance from the transmitting antenna [49, 23]. The strength of the received signal is used to determine the distance between the point of the signal and the point of reception. The calculation formula for RSS is shown in Equation 1 [54]:
@@ 3.4 System Architecture
The system utilises a MySQL database for the purpose of data storage. The data to be stored comprises the location coordinates (X and Y) and the signal strength of the Wi-Fi deployment point [53, 28]. The data types are of the float variety, and none of them permit null values. The design is illustrated in Table 1.
@@ 5 Conclusion
This paper preliminarily explores the classification of indoor position technology and the implementation scheme based on location fingerprint technology. Common indoor localization techniques include computer vision-based localization, sensor-based localization, and localization based on different communication technologies such as RFID, Wi-Fi, Bluetooth, visible light communication, and ultra-wideband. Location fingerprinting has emerged as a prominent technique in this field, using widely deployed Wi-Fi networks. This study proposes a pragmatic hybrid indoor positioning system that integrates Wi-Fi fingerprinting and pedestrian dead reckoning, providing a cost-effective solution that optimizes accuracy, real-time performance, and deployment complexity [30, 36]. By using existing Wi-Fi infrastructure and smartphone sensors, the system can perform data acquisition, localization, and trajectory tracking functions without additional hardware [47, 55]. The proposed dynamic weighted WKNN algorithm enhances traditional fingerprint matching by incorporating distance-based weighting, while sensor fusion with PDR effectively achieves real-time trajectory tracking [63]. The experimental results obtained from field tests have confirmed the feasibility of the system, with stable localization accuracy and low latency.